OpenStudy (anonymous):

Find the distance between the two points. Round to the nearest tenth if necessary. (10, –8), (4, 0)

hero (hero):

@jguillory4, what are your thoughts regarding this problem? Do you have an approach to attempt solving this yet?

OpenStudy (anonymous):

Use the distance formula for the two points

OpenStudy (anonymous):

I have an idea how to solve it

hero (hero):

What idea is that @jguillory4? Mind sharing?

Mehek (mehek14):

\(D=\sqrt{(0-(-8)^2+(4-10)^2}\\ D = \sqrt{(8)^2+(-6)^2}\\D=\sqrt{64+36}\)

Mehek (mehek14):

add 64 + 36 = 100 \(\sqrt{100}=10\)

OpenStudy (campbell_st):

why not find the distance between the x values, then the distance between the y values and just use Pythagoras' Theorem... the distance between the points is the hypotenuse it means less to remember

hero (hero):

@alexxmorales the point was to figure out what the user already knows first before attempting to help.

OpenStudy (campbell_st):



OpenStudy (campbell_st):

so always ask if they can plot the points. then if the can find the distance, then use pythagoras